# Explaining Grokking and Information Bottleneck through Neural Collapse Emergence

### v1

@@ C.1 Grokking
To validate our findings across diverse datasets, we conducted experiments on three text classification benchmarks: SST-2, a binary sentiment classification task (Socher et al., 2013); TREC-6, a question classification task with six classes (Hovy et al., 2001; Li & Roth, 2002); and AG-news, a topic classification task with four news categories (Zhang et al., 2015). For preprocessing, we use the Hugging Face Bert WordPiece tokenizer (Devlin et al., 2019) solely for tokenization, restricting the vocabulary to tokens present in the training set. All embeddings, including unknown and padding tokens, are randomly initialized, and the maximum sequence length is set to $128$. The model configuration is the same as in the MNIST ViT experiment. Training is performed with a weight decay of $1\mathrm{e}{-2}$ and a training set size of $3000$. We show results for both standard initialization and a variant where the initialization scale of the feedforward and linear layers is scaled up by a factor of $8$, building on the prior grokking studies.

### v2

@@ D.1 Grokking
To validate our findings across diverse datasets, we conducted experiments on three text classification benchmarks: SST-2, a binary sentiment classification task (Socher et al., 2013); TREC-6, a question classification task with six classes (Hovy et al., 2001; Li and Roth, 2002); and AG-news, a topic classification task with four news categories (Zhang et al., 2015). For preprocessing, we use the Hugging Face Bert WordPiece tokenizer (Devlin et al., 2019) solely for tokenization, restricting the vocabulary to tokens present in the training set. All embeddings, including unknown and padding tokens, are randomly initialized, and the maximum sequence length is set to $128$. The model configuration is the same as in the MNIST ViT experiment. Training is performed with a weight decay of $1\mathrm{e}{-2}$ and a training set size of $3000$. We show results for both standard initialization and a variant where the initialization scale of the feedforward and linear layers is scaled up by a factor of $8$, building on the prior grokking studies.